Grade 11Mechanics

A 93 kg man lower himself to the ground from a height 10.0 m by holding onto a rope that runs over a frictionless pulley to a 65 kg sandbag with what speed does the man hit the ground if he started from rest?

Harsh6
8 Years agoGrade 11

1 Answer

Devansh Kaushik
8 Years ago
The answer shall be 5.953 ~ 6 m/s. As per the question. The 95kg man hold downs the rope by which he covers a distance of 10m with initially at rest. The rope passes by a pulley to a 65 kg sandbag. So tension acts on the rope. By forming the equations we get93g - T = 93a andT - 65g = 65a where a is the acceleration of the the man. By solving the equation we get a = 1.77 m/s. Now using v2 - u2 = 2as, we get v2 - 02 = 2 x 1.77 x 10 => v2 = 35.4 => v = √(35.44) => v =5.953 ~ 6 m/s